The reflecting surfaces of two intersecting flat m

Calculate the angle ?. Answer in units of ? .
Answer #1

Angle between incident ray and mirror 1 = 53
Angle between reflected ray and mirrir 1 = 53 ( i = r)
using angle sum property in triangle
Angle between incident ray and mirror = 180 - 53 - 53 = 74 deg
Angle between mirror 2 and reflected ray = 74 deg (i = r)
Angle between hortizontal and final ray = 53 - 74 = -21 deg
Angle between incident ray and horizontal = 53 deg
Net angle phi = 53 + 21 = 74 deg